OpenStudy (anonymous):

How to symplify this radical expression? -√49 (thats supposed to be a square root sign)

OpenStudy (kainui):

Well, what's the square root of 49? Is there any number you can think of that you can multiply by itself to get 49?

OpenStudy (anonymous):

I know the square root of 49 is 7 I'm just confused about the negative sign

OpenStudy (anonymous):

Do I have to use i

OpenStudy (anonymous):

You only have to use i if the negative sign is inside the square root symbol

OpenStudy (kainui):

Nope, you only have to use i if the negative sign is inside the square root. Because if you think about it, (-2)^2 is 4 and 2^2 is 4, so you will never get a negative number if you multiply two negatives together. i=square root(-1)

OpenStudy (anonymous):

so then what would I do is it was say \[\sqrt{-81}\]

OpenStudy (kainui):

\[\sqrt{-81}=\sqrt{81*-1}=9i\]

OpenStudy (anonymous):

okay thanks